Bruce has $5.50 in dimes and quarters. He has 8 more quarters then dimes. How many coins does he have?
Found 2 solutions by harpazo, algebrahouse.com:
Answer by harpazo(655)


I will set up the system of equations. You can solve using the substitution method.
10d + 25q = 550...Equation 1
q = 8 + n...Equation 2

Answer by algebrahouse.com(1659)

x = number of dimes
x + 8 = number of quarters {he has 8 more quarters than dimes}

10x + 25(x + 8) = 550 {value of coin times number of coins equals total value}
10x + 25x + 200 = 550 {used distributive property}
35x + 200 = 550 {combined like terms}
35x = 350 {subtracted 200 from each side}
x = 10 {divided each side by 35}
x + 8 = 18 {substituted 8, in for x, into x + 8}

10 dimes and 18 quarters

For more help from me, visit: www.algebrahouse.com